Let X be a random variable with mean 6 and standard deviation 4. If L = 8-3X what is the standard deviation of L?
Answer #1

E(X ) = mean for x = 6

We calculate the variance of x by using the standard deviation of

Standard deviation (x) = 4

Var(X) = [\ standard \ deviation \ (x)\ ]^2

Var(X) = [4]^2 = 16

Formula

Var( a - b*X) = b2 * Var(X)

Var(L) = Var(8-3x) = 3^2 *Var(X)= 9*16 = 144

Var(L) = 144

Now we find a standard deviation of L

Standard \ deviation \ of \ L = \sqrt{var(L)}

Standard \ deviation \ of \ L = \sqrt{144}

Standard \ deviation \ of \ L ={\color{Purple} \mathbf{12}}

Another way

Standard deiation ( a - b*x ) = b * standard deviation (x)

Standard deviation of (L ) = standard deviation ( 8 - 3*x)

Standard deviation of (L ) = 3* Standard deviation (x)

We have standard deviation of (x) = 4

Standard deviation of (L ) = 3*4 = 12

Final answer:-

Standard deviation of (L ) = 12